CW FIBER LASER CAVITY RING DOWN SPECTROSCOPY

\underline{WEIDONG CHEN}, JULIEN COUSIN, Laboratoire de Physicochimie de l’Atmosphère, Université du Littoral Côte d’Opale, 189A, Av. Maurice Schumann, 59140 Dunkerque, France; SAMIR KASSI, ALAIN CAMPARGUE, DANIELE ROMANINI, Laboratoire de Spectrométrie Physique, Université J. Fourier de Grenoble, 140 rue de la Physique, 38402 St Martin d’Hères, France.

Application of a tunable near-infrared Erbium doped fiber laser to continuous-wave (cw) cavity ring down spectroscopy (CRDS) is reported. The laser is widely tunable in the spectral region of 1542-1600 nm by temperature tuning with a laser linewidth of ∼ kHz. This ultranarrow linewidth is highly desirable for efficient cw laser coupling to a high-finesse ring down cavity^{a}. Cavity ring down spectra of N_{2}O and C_{2}H_{2} have been recorded near 1.544 μm. Measurements of N_{2}O concentration have been performed as well with high dynamic range varying from 100 % purity to some tens ppb with a minimum measurable absorption coefficient of about 10^{−10} cm^{−1}. Preliminary results will be presented.

aJ. Morville, D. Romanini, M. Chenevier, A. Kachanov, Appl. Opt. \underline{41}, 6980 (2002).